OpenStudy (anonymous):

find vertical and horizontal asymptotes of a curve: y=(x^(2)+1)/(5x^(2)-44x-9)...please show steps thanks in advance: )

OpenStudy (anonymous):

is it really \[y=\frac{x^2+1}{5x^2-44x-9}\]

OpenStudy (anonymous):

yes satellite73

OpenStudy (anonymous):

vertical asymptote: set denominator = 0 and solve

OpenStudy (anonymous):

by some miracle this actually factors as \[(5x+1)(x-9)=0\] so vertical asymptotes are \[x=-\frac{1}{5}\] and \[x=9\]

OpenStudy (anonymous):

ok so would be lik 5x^2-44x-9=0 and it would be like this?

OpenStudy (anonymous):

horizontal asymptote much easier. the degrees are the same top and bottom (both 2) so take \[y=\frac{\text{leading coefficient}}{\text{leading coefficient}}\]

OpenStudy (anonymous):

ok so you factor it out right?

OpenStudy (anonymous):

1/9?

OpenStudy (anonymous):

yes that is right. put \[5x^2-44x-9=0\] and solve for x

OpenStudy (anonymous):

i mean 1/5

OpenStudy (anonymous):

nope you factored wrong

OpenStudy (anonymous):

still wrong. it is -1/5 or 9

OpenStudy (anonymous):

wait for horizontal?

OpenStudy (anonymous):

no no those are vertical. vertical where denominator is zero

OpenStudy (anonymous):

the leading coeffeicent would be 1/5

OpenStudy (anonymous):

the ratio of leading coefficients is 1/5 yes. so that is horizontal asymptote. you got it

OpenStudy (anonymous):

yes but i meant to give answer for the horizontal section

OpenStudy (anonymous):

yes it is \[y=\frac{1}{5}\]

OpenStudy (anonymous):

yes thanks soo much!! : )

OpenStudy (anonymous):

yw